### v1

@@
9
The hard-core model in graph theory [^1] This survey has been prepared as a chapter for a forthcoming volume, Topics in Probabilistic Graph Theory. Date of draft: 6 January 2025.
Ewan Davies [^2] Department of Computer Science, Colorado State University, Fort Collins, USA. Supported in part by NSF grant CCF-2309707. research@ewandavies.org   and   Ross J. Kang [^3] Korteweg–de Vries Institute for Mathematics, University of Amsterdam, Netherlands. Supported by Vidi (639.032.614) and Open Competition (OCENW.M20.009) grants of the Dutch Research Council (NWO) as well as the Gravitation programme NETWORKS (024.002.003) of the Dutch Ministry of Education, Culture and Science (OCW). ross.kang@gmail.com
@@ 1 Introduction
In this chapter, our aim is not to give the most comprehensive or in-depth treatment of the use of the hard-core model in graph theory, which would indeed be a bit hard-core (and perhaps self-defeating), but rather to give a current overview of some of its most elegant and interesting uses, and to provide a glimpse into its varied applicability. In this section, with the help of a ‘toy’ example, we introduce the model. Then in the following section, we sketch a series of ‘vignettes’, where the hard-core model, and in particular the local occupancy method, has had influence. In later sections, we dive into more of the details.
Given a graph, consider a uniformly random independent set $X$. To be completely clear, we first gather all of the independent sets of the graph (including $\varnothing$) into a collection $\mathcal{I}$, and then choose $X$ randomly from $\mathcal{I}$, where each member of $\mathcal{I}$ is chosen with equal probability. Let $u$ be a vertex and reveal the number $Y_{u}=|X\cap N(u)|$ of neighbours of $u$ occupied by $X$. Since $X$ is an independent set, the event $u\in X$, that $u$ itself is occupied, cannot occur unless $Y_{u}=0$. Moreover, since $X$ is chosen uniformly from $\mathcal{I}$, it is intuitive that, conditional on $Y_{u}=0$, the events $u\in X$ and $u\notin X$ have equal probability (and we prove this rigorously in more generality below). Thus, $\mathbb{P}(u\in X)=\frac{1}{2}\mathbb{P}(Y_{u}=0)=q/2$, where we write $q=\mathbb{P}(Y_{u}=0)$. Markov’s inequality then implies that $\mathbb{E}|X\cap N(u)|=\mathbb{E}Y_{u}\geq\mathbb{P}(Y_{u}\geq 1)=1-q$. It follows that
$12\mathbb{P}(u\in X)+\mathbb{E}|X\cap N(u)|\geq 2\cdot\tfrac{1}{2}q+(1-q)=$ 1 .
@@
and so the only information about $I$ on which this probability depends is $|I|$. We have recovered the hard-core model on $G$ at fugacity $p/(1-p)$.
@@
A little calculus shows then that $p_{I}=\exp(-1+\Lambda_{1}\log 2+|I|\Lambda_{2}\log 2)$, and again there must be some value of $\lambda$ such that $p_{I}\propto\lambda^{|I|}$. The precise value of $\lambda$ is a function of $\mu$, and for the rest of the calculation, we require that $0\leq\mu\leq\alpha(G)$. These properties mean that the hard-core model arises naturally in physics as a discrete model of gas particles that are constrained to the vertices of $G$. The only interaction that these particles have is that they cannot simultaneously occupy both endpoints of an edge of $G$, and this self-exclusion property forms the basis of a discrete model of a gas (the so-called hard-core lattice gas) where particles have fixed size and cannot overlap, but are not subject to inter-particle forces. For more of a physics perspective, see, for example, [70][36], and [38].
@@ 2 Vignettes
It is simple, but crucial, to notice that in any triangle-free graph, the set of neighbours of any given vertex forms an independent set. It is therefore intuitive that triangle-freeness might lead to there being larger independent sets than in general, as there are independent sets spread locally throughout the graph! Indeed, if we are interested in finding an independent set of size $k$ for any triangle-free graph on $n$ vertices, we may at least assume without loss of generality that all vertices in the graph have degree less than $k$. We have thereby reduced the problem of showing $R(3,k)\leq n$ to that of finding an independent set of size $k$ in any triangle-free graph on $n$ vertices of maximum degree at most $k-1$. It turns out that the more general problem of finding large independent sets in triangle-free graphs on $n$ vertices of maximum degree at most $\Delta$ is the crux of bounding $R(3,k)$ from above, especially asymptotically, as $k\to\infty$.
Question: In a triangle-free graph on $n$ vertices of maximum degree at most $\Delta$, how large an independent set can we guarantee?
The seminal work of Ajtai, Komlós and Szemerédi [2][3] – which incidentally also introduced an early form of the nibble method – showed in this setting that the independence number must be at least $\Omega((n\log\Delta)/\Delta)$, improving on the ‘trivial’ bound presented in the Introduction by a factor logarithmic in $\Delta$. Not long afterwards, Shearer [65] improved this independence number bound by a constant factor to $(1+o(1))(n\log\Delta)/\Delta$ as $\Delta\to\infty$. (Unless mentioned otherwise, the base of the logarithm is always natural.) If we substitute $\Delta=k-1$ as in the previous paragraph, then we should clearly take $n=(1+o(1))k^{2}/\log k$ to derive that $R(3,k)\lesssim k^{2}/\log k$. Despite sustained attention from many leading researchers, Shearer’s bound remains asymptotically the state-of-the-art upper bound for $R(3,k)$. The earlier-mentioned work on the triangle-free process (see [14][35]) shows this to be counterbalanced by a lower bound of the form $R(3,k)\gtrsim k^{2}/(4\log k)$.
@@
Question: Given a collection of non-intersecting congruent spheres in $\mathbb{R}^{d}$, how high can their density in $\mathbb{R}^{d}$ be?
@@
Inspired by the corresponding finite results, Jenssen, Joos and Perkins [45] studied directly the analogue of the hard-core model on the infinite graph on $\mathbb{R}^{d}$, known as the hard-sphere model. They proved that a random packing from this model achieves density $\Omega(d2^{-d})$ in expectation, and their techniques evoked the concept of local occupancy that we study here in the continuous setting. The best-known lower bound $\theta(d)\geq\Omega(d2^{-d}\log d)$ is due to Campos, Jenssen, Michelen and Sahasrabudhe [21] and relies on a novel disordered discretization of space, as well as a new lower bound on the independence number of locally sparse graphs. Apart from determining $\theta(d)$, the question of whether dense packings in high dimensions are disordered or lattice-like remains elusive.
@@
Question: Given a triangle-free graph of maximum degree $\Delta$, how large can the chromatic number be?
In general, this problem is very difficult. Even the case $\Delta=5$ remains open: in 1970, Grünbaum [39] conjectured that there exists some triangle-free $5$-regular graph of chromatic number $5$. While others conjectured the opposite (see [60]), it remains open to debate. For other small values of $\Delta$, an upper bound of $2\lceil(\Delta+2)/3\rceil$ is due to Kostochka in 1982 (see [44, Sec. 4.6]).
In a remarkable unpublished work in the 1990s, using the nibble method (see Chapter 8), Johansson [46] (see also [57]) showed how in triangle-free graphs the asymptotic extremal behaviour, for large $\Delta$, of the chromatic number is aligned to that of the independence number, and showed that the chromatic number may be no larger than $O(\Delta/\log\Delta)$. This implies the Ajtai–Komlós–Szemerédi result by taking a largest colour class in an optimal proper colouring.
@@
When $|\Gamma|=n$, there are at most $2^{o(n)}$ subsets of $\Gamma$ of size less than $\log_{2}n$, and so it suffices to obtain the bound $2^{(1/2+o(1))n}$ on the number of sum-free subsets of $\Gamma$ of size at least $\log_{2}n$. For each such sum-free subset $A$ of $\Gamma$, let $S_{A}$ be the first $\log_{2}n$ elements of $A$. Then $A$ is an $S_{A}$-free subset of $\Gamma$, and so for each fixed $S$ we want an upper bound on the number of independent sets in $G_{S}$, which is regular of degree $|S\cup-S|\geq\log_{2}n$. Supposing that we have established Granville’s conjectured bound, we have that each $S$ can arise as the initial segment of at most $2^{(1/2+o(1))n}$ sum-free sets $A$. The number of possible sets $S$ is $2^{o(n)}$, giving the bound $2^{(1/2+o(1))n}$ on the number of large sum-free sets, as required.
Question: Given a $d$-regular graph on $n$ vertices, how large can the number of independent sets be?
Several solutions to Cameron’s problem appeared around 1990 (see [18, 5] and an unpublished proof due to Erdős and Granville [20]). But it was Alon [5] who tackled Granville’s problem on independent sets in regular graphs. Alon suggested that a sharper bound is plausible: that when $n$ is divisible by $2d$, the tight upper bound, which is attained by a disjoint union of $n/(2d)$ copies of the complete bipartite graph $K_{d,d}$, should be $(2^{d+1}-1)^{n/(2d)}$. With a different motivation, related to statistical physics and phase transitions of the hard-core model, Kahn [47] developed the so-called ‘entropy method’ and established this bound in the case of bipartite regular graphs. Later, Zhao [74] reduced the general case to the bipartite case, thereby settling the problem for all regular graphs. These results were extended to bounds on the partition function of the hard-core model and to more general weighted counting problems by Kahn [48] and Galvin and Tetali [37].
For the hard-core model, the results above were strengthened to bounds on the expected size of the independent set by Davies, Jenssen, Perkins and Roberts [26]. Their method readily incorporated local structure, which offers advantages in some settings. Their methods could be applied to triangle-free graphs (see [27]) to produce independence number bounds analogous to Shearer’s. In fact, Shearer knew that the hard-core model could be used in this way and presented his work at the 1998 SIAM annual meeting and conference on discrete mathematics held in Toronto [68]. One of the open problems Shearer highlighted was the question of whether the analysis of the hard-core model could provide an efficient algorithm for constructing independent sets of the size certified by the method and, through the work of Molloy [56] and Bonamy, Kelly, Nelson and Postle [15], these threads were finally united in the form of local occupancy (see [30]).
@@ 3 Random independent sets: a balancing act
Our objective in this section is to motivate how various interesting local properties of the hard-core model at fugacity $\lambda$ lead naturally to nice bounds on the occupancy fraction, the independence number and the fractional chromatic number, subject to an analytic optimization in terms of $\lambda$ and the maximum degree. We remark that, while the power to optimize over $\lambda$ is handy, the specialization to the uniform case $\lambda=1$ already contains much of the important intuition. We have given a hint of this line of reasoning in the Introduction, but we now provide extra detail and intuition.
@@
Many of the methods we describe here take advantage of an elegant Markovian property of the hard-core model, which allows us to isolate our considerations between different parts of the graph. Suppose that we want to know how $X$ is distributed with respect to the subgraph $G[A]$ induced by some subset $A\subseteq V$. Of course, if we expose $X$ outside of $A$, then we must exclude consideration of those vertices in $A$ that have some neighbour (external to $A$) occupied by $X$. The spatial Markov property asserts that it is sufficient to independently instantiate the distribution on those vertices in $A$ that remain, the so-called externally uncovered vertices in $A$. That assertion is formalized in the following result.
@@
Proof.
Let $I_{0}$ be an arbitrary independent set of $G[V\setminus A]$, and let us condition on the fact that $X\setminus A=I_{0}$. It follows that $X\cap A$ is a subset of $A\setminus N(I_{0})$, and indeed is an independent set in $G[A\setminus N(I_{0})]=F_{A}$. For any independent set $I_{1}$ in $F_{A}$, we have
@@
and this completes the proof because $Z_{F_{A}}(\lambda)=\sum_{J\in\mathcal{I}(F_{A})}\lambda^{|J|}$. ∎
@@
$\mathbb{P}(X\cap A=\varnothing)=\mathbb{E}\frac{1}{Z_{F_{A}}(\lambda)}\quad\text{ and }\quad\mathbb{E}|X\cap A|=\mathbb{E}\frac{\lambda Z_{F_{A}}^{\prime}(\lambda)}{Z_{F_{A}}(\lambda)}.$
Proof.
We let $A_{0}$ be an arbitrary subset of $A$, and we condition on the event that $A\setminus N(X\setminus A)=A_{0}$. By Theorem 3.1, we have
@@
which is as claimed in the Introduction. It follows from Corollary 3.2 with $A=N(u)$ that
@@
Therefore, what Corollary 3.2 suggests is that in order to establish conditions of the form in (9.1), it is useful to understand the local behaviour of the hard-core model. In particular, we want to understand $Z_{F}(\lambda)$ for induced neighbourhood subgraphs $F$ in $G$, that is, for the induced subgraphs $F$ of $G[N(u)]$ for $u\in V$. These remarks motivate the following definition, which is central to the chapter.
Definition 3.3.
@@
In the next section, we give some examples of how we can find more tailored choices for $(\beta,\gamma)$ than the ones in the Introduction, given various local structural assumptions on $G$.
@@
If $X$ has local $(\beta,\gamma)$-occupancy as per Definition 3.3, then the following hold.
@@
Proof.
For part (a), we choose the distribution of $X_{B}$ as the hard-core model on $G[B]$ at fugacity $\lambda$. Since Corollary 3.2 and the accompanying remarks also apply to $X_{B}$ by our assumption on $X$, we derive the inequality immediately from the condition (9.4) with $F$ chosen as $G[N_{G[B]}(u)]$, which is an induced subgraph of $G[N(u)]$.
Part (b) follows similarly from Corollary 3.2 and the accompanying remarks, where we take $F$ to be the entire neighbourhood subgraph $G[N(u)]$, and then average over the choice of $u$. ∎
Of the conditions of Definition 3.3, Theorem 3.4(a) and Theorem 3.4(b), the first is the strongest while the third is in a sense the weakest. However, in practice our local optimization analyses in the hard-core model do not yield any different $\beta,\gamma$ when we target these conditions separately. In fact, it is the resemblance between these conditions that lies at the root of this chapter. Thus, in an obvious abuse, we may refer to any of these three conditions as verifying local $(\beta,\gamma)$-occupancy for the distribution of $X$.
The reason, however, for making explicit these three versions of local occupancy is that they lead to three different implications for the global (independent set) structure of $G$. Moreover, whereas we have assumed throughout the section that $X$ has the hard-core distribution, the assertions in parts (a) and (b), respectively, of Theorem 3.4 are the essential hypotheses for the conclusions of the following two results. In other words, for these results, we do not in principle need that $X$ be distributed according to the hard-core model.
@@
If $X$ satisfies local $(\beta,\gamma)$-occupancy for some $\beta,\gamma>0$ in the sense of Theorem 3.4, part (b), then
@@
Proof.
By our assumption of the condition in Theorem 3.4, part (b), the sum of
@@
where the inequality holds because each vertex $v$ appears at most $\Delta$ times in the double sum. Rearranging then yields the desired result. ∎
While the occupancy fraction bound is a simple application of local occupancy, it is a crucial aspect of the technique, because in some settings we can prove that the given bound is asymptotically tight. This is not necessarily true for applications to the independence number or to the (fractional) chromatic number. Another important point is that, since larger values of $\lambda$s bias to larger independent sets, the expectation $\mathbb{E}|X|$ is monotone increasing in $\lambda$ for any fixed graph (see [27, Prop. 1]), but the quantity $1/(\beta+\gamma\Delta)$ with which we bound it might not be. One reason that the case of triangle-free graphs is more subtle and challenging than the general case described in the Introduction is precisely this behaviour of $\beta+\gamma\Delta$ as $\lambda$ varies.
@@
Although there are several equivalent definitions of fractional colouring, we prefer a probabilistic one, in terms of uniform occupancy. A fractional $k$-colouring of $G$ is a distribution over $\mathcal{I}$ for which it holds, writing $X_{f}$ for the random independent set, that $\mathbb{P}(u\in X_{f})\geq 1/k$ for each $u\in V$. Note that by summing this over all $u\in V$, we have $\mathbb{E}|X_{f}|\geq|V|/k$. The fractional chromatic number $\chi_{f}(G)$ is the least $k$ for which $G$ admits a fractional $k$-colouring. Note also that one can see fractional colouring as a relaxation of usual proper colouring, with $\chi_{f}(G)\leq\chi(G)$ always, by noting the simple fractional $\chi(G)$-colouring that selects uniformly from one of the $\chi(G)$ independent sets induced by an optimal proper colouring of $G$.
@@
If $X$ satisfies local $(\beta,\gamma)$-occupancy for some $\beta,\gamma>0$ in the sense of Theorem 3.4, part (a), then
@@
Proof.
Let us write $k$ for the desired bound $\beta+\gamma\Delta$. We proceed to build up a fractional colouring ${w}$ in several iterations, and we write ${w}(I)$ for the probability mass assigned to the independent set $I$ and ${w}(u)$ for the total probability mass $\sum_{I\in\mathcal{I},I\ni u}{w}(I)$ assigned to independent sets containing the vertex $u$. Through the iterations, ${w}$ is a partial fractional $k$-colouring of $G$, in the sense of not yet having satisfied the condition that ${w}(u)\geq 1/k$ for all $u\in V$, while the total weight ${w}(G)=\sum_{I\in\mathcal{I}}{w}(I)$ over $\mathcal{I}$ remains at most $1$. The idea is to accrue probability mass to the independent sets according to $X_{V(H)}$, for the relevant induced subgraph $H$ of $G$, until the mass containing some vertex is $1/k$, whereupon we remove that vertex from $H$ and continue. This idea is made precise in Algorithm 1.

### v2

@@
8
The hard-core model in graph theory  [^1] This survey has been prepared as a chapter for a forthcoming volume, Topics in Probabilistic Graph Theory. Draft dated 3 September 2025.
Ewan Davies  [^2] Department of Computer Science, Colorado State University, Fort Collins, USA. research@ewandavies.org   and   Ross J. Kang  [^3] Korteweg–de Vries Institute for Mathematics, University of Amsterdam, Netherlands. ross.kang@gmail.com
[…]
@@ 1 Introduction
In this chapter, our aim is not to give the most comprehensive or in-depth treatment of the use of the hard-core model in graph theory, which would indeed be a bit hard-core (and perhaps self-defeating), but rather to give a current overview of some of its most elegant and interesting uses, and to provide a glimpse into its varied applicability. In this section, with the help of a ‘toy’ example, we introduce the model. Then in the following section, we sketch a series of ‘vignettes’, where the hard-core model, and in particular the local occupancy method, has had influence. In later sections, we dive into more of the details of a recent framework for inferring graph structure that was developed around the hard-core model. At the end, we discuss some frontiers related to the methods and topics covered in this chapter.
A toy example
Given a graph, consider a uniformly random independent set $X$. To be completely clear, we first gather all of the independent sets of the graph (including $\varnothing$) into a collection $\mathcal{I}$, and then choose $X$ randomly from $\mathcal{I}$, where each member of $\mathcal{I}$ is chosen with equal probability. Let $u$ be a vertex and reveal the number $Y_{u}=|X\cap N(u)|$ of neighbours of $u$ occupied by $X$. Since $X$ is an independent set, the event $u\in X$, that $u$ itself is occupied, cannot occur unless $Y_{u}=0$. Moreover, since $X$ is chosen uniformly from $\mathcal{I}$, it is intuitive that, conditional on $Y_{u}=0$, the events $u\in X$ and $u\notin X$ have equal probability (and we prove this rigorously in more generality below). Thus, $\mathbb{P}(u\in X)=\frac{1}{2}\mathbb{P}(Y_{u}=0)=\tfrac{1}{2}q$, where we write $q=\mathbb{P}(Y_{u}=0)$. Markov’s inequality then implies that $\mathbb{E}|X\cap N(u)|=\mathbb{E}Y_{u}\geq\mathbb{P}(Y_{u}\geq 1)=1-q$. It follows that
$2\mathbb{P}(u\in X)+\mathbb{E}|X\cap N(u)|\geq 2\cdot\tfrac{1}{2}q+(1-q)=1.$
@@
and so the only information about $I$ on which this probability depends is $|I|$. We have recovered the hard-core model on $G$ at fugacity p/(1p)p/(1-p).
@@
A little calculus shows then that $p_{I}=\exp(-1+\Lambda_{1}\log 2+|I|\Lambda_{2}\log 2)$, and again there must be some value of $\lambda$ such that $p_{I}\propto\lambda^{|I|}$. (Unless specified otherwise, the base of the logarithm is always natural.) The precise value of $\lambda$ is a function of $\mu$, and for the rest of the calculation, we require that $0\leq\mu\leq\alpha(G)$. These properties mean that the hard-core model arises naturally in physics as a discrete model of gas particles that are constrained to the vertices of $G$. The only interaction that these particles have is that they cannot simultaneously occupy both endpoints of an edge of $G$, and this self-exclusion property forms the basis of a discrete model of a gas (the so-called hard-core lattice gas) where particles have fixed size and cannot overlap, but are not subject to inter-particle forces. For more of a physics perspective, see, for example, [11][41], and [43].
@@ 2 Vignettes
It is simple, but crucial, to notice that in any triangle-free graph, the set of neighbours of any given vertex forms an independent set. It is therefore intuitive that triangle-freeness might lead to there being larger independent sets than in general, as there are independent sets spread locally throughout the graph! Indeed, if we are interested in finding an independent set of size $k$ for any triangle-free graph on $n$ vertices, we may at least assume without loss of generality that all vertices in the graph have degree less than $k$. We have thereby reduced the problem of showing $R(3,k)\leq n$ to that of finding an independent set of size $k$ in any triangle-free graph on $n$ vertices of maximum degree at most k1k-1. It turns out that the more general problem of finding large independent sets in triangle-free graphs on $n$ vertices of maximum degree at most $\Delta$ is the crux of bounding $R(3,k)$ from above, especially asymptotically, as $k\to\infty$.
Question
In a triangle-free graph on $n$ vertices of maximum degree at most $\Delta$, how large an independent set can we guarantee?
The seminal work of Ajtai, Komlós and Szemerédi [3][4] – which simultaneously introduced an early form of the nibble method – showed in this setting that the independence number must be at least $\Omega((n\log\Delta)/\Delta)$, improving on the ‘trivial’ bound presented in Section 1 by a factor logarithmic in $\Delta$. Not long afterwards, Shearer [71] improved this independence number bound by a constant factor to $(1+o(1))(n\log\Delta)/\Delta$ as $\Delta\to\infty$. If we substitute $\Delta=k-1$ as in the previous paragraph, then we should clearly take $n=(1+o(1))k^{2}/\log k$ to derive that $R(3,k)\lesssim k^{2}/\log k$. Note that these independence number results apply in the more general setting of given average degree, though one does not need this additional strength for the Ramsey number bound.
Despite sustained attention from many leading researchers, Shearer’s bound remains asymptotically the state-of-the-art upper bound for $R(3,k)$. The earlier-mentioned work on the triangle-free process (see [16][40]) shows this to be counterbalanced by a lower bound of the form $R(3,k)\gtrsim k^{2}/(4\log k)$.
@@
Question
Given a collection of non-intersecting congruent spheres in $\mathbb{R}^{d}$, how high can their density in $\mathbb{R}^{d}$ be?
@@
Inspired by the corresponding finite results, Jenssen, Joos and Perkins [50] studied directly the analogue of the hard-core model on the infinite graph on $\mathbb{R}^{d}$, known as the hard-sphere model. They proved that a random packing from this model achieves a density of $\Omega(d2^{-d})$ in expectation, and their techniques evoked the concept of local occupancy that we study here in the continuous setting. The best-known lower bound $\theta(d)\geq\Omega(d2^{-d}\log d)$ is due to Campos, Jenssen, Michelen and Sahasrabudhe [24] and relies on a novel disordered discretization of space, as well as a new lower bound on the independence number of locally sparse graphs. Apart from determining $\theta(d)$, the question of whether dense packings in high dimensions are disordered or lattice-like remains elusive.
@@
Question
Given a triangle-free graph of maximum degree $\Delta$, how large can the chromatic number be?
In general, this problem is very difficult. Even the case $\Delta=5$ remains open: in 1970, Grünbaum [44] conjectured that there exists some triangle-free $5$-regular graph of chromatic number $5$. While others conjectured the opposite (see [66]), it remains open to debate. For other small values of $\Delta$, an upper bound of $2\lceil(\Delta+2)/3\rceil$ is due to Kostochka in 1982 (see [49, Sec. 4.6]).
In a remarkable unpublished work in the 1990s, using the nibble method (see Chapter 7), Johansson [51] (see also [63]) showed how in triangle-free graphs the asymptotic extremal behaviour, for large $\Delta$, of the chromatic number is aligned to that of the independence number, and showed that the chromatic number may be no larger than $O(\Delta/\log\Delta)$. This implies the Ajtai–Komlós–Szemerédi result by taking a largest colour class in an optimal proper colouring.
@@
When $|\Gamma|=n$, there are at most $2^{o(n)}$ subsets of $\Gamma$ of size less than $\log_{2}n$, and so it suffices to obtain the bound $2^{(1/2+o(1))n}$ on the number of sum-free subsets of $\Gamma$ of size at least $\log_{2}n$. For each such sum-free subset $A$ of $\Gamma$, let $S_{A}$ be the first $\log_{2}n$ elements of $A$. Then $A$ is an $S_{A}$-free subset of $\Gamma$, and so for each fixed $S$ we want an upper bound on the number of independent sets in $G_{S}$, which is regular of degree $|S\cup(-S)|\geq\log_{2}n$. Supposing that we have established Granville’s conjectured bound, we have that each $S$ can arise as the initial segment of at most $2^{(1/2+o(1))n}$ sum-free sets $A$. The number of possible sets $S$ is $2^{o(n)}$, giving the bound $2^{(1/2+o(1))n}$ on the number of large sum-free sets, as required.
Question
Given a $d$-regular graph on $n$ vertices, how large can the number of independent sets be?
Several solutions to Cameron’s problem appeared around 1990 (see [21][6] and an unpublished proof due to Erdős and Granville [23]). But it was Alon [6] who tackled Granville’s problem on independent sets in regular graphs. Alon suggested that a sharper bound is plausible: that when $n$ is divisible by $2d$, the tight upper bound, which is attained by a disjoint union of $n/(2d)$ copies of the complete bipartite graph $K_{d,d}$, should be $(2^{d+1}-1)^{n/(2d)}$. With a different motivation, related to statistical physics and phase transitions of the hard-core model, Kahn [52] developed the so-called ‘entropy method’ and established this bound in the case of bipartite regular graphs. Later, Zhao [79] reduced the general case to the bipartite case, thereby settling the problem for all regular graphs. These results were extended to bounds on the partition function of the hard-core model and to more general weighted counting problems by Kahn [53] and Galvin and Tetali [42].
For the hard-core model, the results above were strengthened to bounds on the expected size of the independent set by Davies, Jenssen, Perkins and Roberts [30]. Their method readily incorporated local structure, which offers advantages in some settings. Their methods could be applied to triangle-free graphs (see [31]) to produce independence number bounds analogous to Shearer’s. In fact, Shearer knew that the hard-core model could be used in this way and presented his work at the 1998 SIAM annual meeting and conference on discrete mathematics held in Toronto [74]. One of the open problems that Shearer highlighted was the question of whether the analysis of the hard-core model could provide an efficient algorithm for constructing independent sets of the size certified by the method and, through the work of Molloy [62] and Bonamy, Kelly, Nelson and Postle [17], these threads were finally united in the form of local occupancy (see [35]).
@@ 3 Random independent sets: a balancing act
Our objective in this section is to motivate how various interesting local properties of the hard-core model at fugacity $\lambda$ lead naturally to nice bounds on the occupancy fraction, the independence number and the fractional chromatic number, subject to an analytic optimization in terms of $\lambda$ and the maximum degree. We remark that, while the power to optimize over $\lambda$ is handy, the specialization to the uniform case $\lambda=1$ already contains much of the important intuition. We have given a hint of this line of reasoning in Section 1, but we now provide extra detail and intuition.
@@
Many of the methods that we describe here take advantage of an elegant Markovian property of the hard-core model, which allows us to isolate our considerations between different parts of the graph. Suppose that we want to know how $X$ is distributed with respect to the subgraph $G[A]$ induced by some subset $A\subseteq V$. Of course, if we expose $X$ outside of $A$, then we must exclude consideration of those vertices in $A$ that have some neighbour (external to $A$) occupied by $X$. The spatial Markov property asserts that it is sufficient to independently instantiate the distribution on those vertices in $A$ that remain, the so-called externally uncovered vertices in $A$. That assertion is formalized in the following result.
@@
To see this, let $I_{0}$ be an arbitrary independent set of $G[V\setminus A]$, and let us condition on the event that $X\setminus A=I_{0}$. It follows that $X\cap A$ is a subset of $A\setminus N(I_{0})$, and indeed is an independent set in $G[A\setminus N(I_{0})]=F_{A}$. For any independent set $I_{1}$ in $F_{A}$, we have
@@
and this completes the proof of Theorem 3.1 because $Z_{F_{A}}(\lambda)=\sum_{J\in\mathcal{I}(F_{A})}\lambda^{|J|}$.
@@
$\mathbb{P}(X\cap A=\varnothing)=\mathbb{E}\frac{1}{Z_{F_{A}}(\lambda)}\quad\text{ and }\quad\mathbb{E}|X\cap A|=\mathbb{E}\frac{\lambda Z_{F_{A}}^{\prime}(\lambda)}{Z_{F_{A}}(\lambda)}.$ (8.4)
For this derivation, we let $A_{0}$ be an arbitrary subset of $A$, and we condition on the event that $A\setminus N(X\setminus A)=A_{0}$. By Theorem 3.1, we have
@@
which is as claimed in Section 1. It follows from (8.4) with $A=N(u)$ that
@@
Therefore, what equation (8.4) suggests is that in order to establish conditions of the form in (8.1), it is useful to understand the local behaviour of the hard-core model. In particular, we want to understand $Z_{F}(\lambda)$ for induced neighbourhood subgraphs $F$ in $G$, that is, for the induced subgraphs $F$ of $G[N(u)]$ for $u\in V$. These remarks motivate the following definition, which is central to the chapter.
Key Definition
@@
In the next section, we give some examples of how we can find more tailored choices for $(\beta,\gamma)$ than the ones in Section 1, given various local structural assumptions on $G$.
@@
If $X$ has local $(\beta,\gamma)$-occupancy, then the following hold.
@@
For part (a), we choose the distribution of $X_{B}$ as the hard-core model on $G[B]$ at fugacity $\lambda$. With $F$ chosen as $G[N_{G[B]}(u)]$, it is an induced subgraph of $G[N(u)]$. Since (8.4) and the accompanying remarks also apply to $X_{B}$ by our assumption on $X$, we derive the inequality immediately from condition (8.5). Part (b) follows similarly from (8.4) and the accompanying remarks, where we take $F$ to be the entire neighbourhood subgraph $G[N(u)]$, and then average over the choice of $u$.
Of the conditions of the Key Definition, Theorem 3.2(a) and Theorem 3.2(b), the first is the strongest while the third is in a sense the weakest. However, in practice our local optimization analyses in the hard-core model do not yield any different values for $\beta$ and $\gamma$ when we target these conditions separately. In fact, it is the resemblance between these conditions that lies at the root of this chapter. Thus, in an obvious abuse, we may refer to any of these three conditions as verifying local $(\beta,\gamma)$-occupancy for the distribution of $X$.
The reason, however, for making explicit these three versions of local occupancy is that they lead to three different implications for the global (independent set) structure of $G$. Moreover, whereas we have assumed throughout the section that $X$ has the hard-core distribution, the assertions in parts (a) and (b) of Theorem 3.2 are the essential hypotheses for the conclusions of the following two results. In other words, for these results, we do not in principle need that $X$ be distributed according to the hard-core model.
@@
If $X$ satisfies local $(\beta,\gamma)$-occupancy for some $\beta,\gamma>0$ in the sense of Theorem 3.2(b), then
@@
To see this result, note that, by the condition in Theorem 3.2(b), the sum of
@@
where the inequality holds because each vertex $v$ appears at most $\Delta$ times in the double sum. Rearranging then yields the desired result.
[…]
@@
Although there are several equivalent definitions of fractional colouring, we prefer a probabilistic one, in terms of uniform occupancy. A fractional $k$-colouring of $G$ is a distribution over $\mathcal{I}$ for which it holds that $\mathbb{P}(u\in X_{f})\geq 1/k$ for each $u\in V$, where $X_{f}$ denotes the random independent set. Note that by summing this over all $u\in V$, we have $\mathbb{E}|X_{f}|\geq|V|/k$. The fractional chromatic number $\chi_{f}(G)$ is the least $k$ for which $G$ admits a fractional $k$-colouring. Note also that one can see fractional colouring as a relaxation of usual proper colouring, with $\chi_{f}(G)\leq\chi(G)$ always, by noting the simple fractional $\chi(G)$-colouring that selects uniformly from one of the $\chi(G)$ independent sets induced by an optimal proper colouring of $G$.
@@
If $X$ satisfies local $(\beta,\gamma)$-occupancy for some $\beta,\gamma>0$ in the sense of Theorem 3.2(a), then
@@
To see how this result holds, let us write $k$ for the desired bound $\beta+\gamma\Delta$. We proceed to build up a fractional colouring w{w} in several iterations, and we write w(I){w}(I) for the probability mass assigned to the independent set $I$ and w(u){w}(u) for the total probability mass $\sum_{I\in\mathcal{I},I\ni u}{w}(I)$ assigned to independent sets containing the vertex $u$. Through the iterations, w{w} is a partial fractional $k$-colouring of $G$, in the sense of not yet having satisfied the condition that ${w}(u)\geq 1/k$ for all $u\in V$, while the total weight ${w}(G)=\sum_{I\in\mathcal{I}}{w}(I)$ over $\mathcal{I}$ remains at most $1$. The idea is to accrue probability mass to the independent sets according to $X_{V(H)}$, for the relevant induced subgraph $H$ of $G$, until the mass containing some vertex is $1/k$, whereupon we remove that vertex from $H$ and continue. This idea is made precise in Algorithm 1.